# MuJo: Multimodal Joint Feature Space Learning for Human Activity Recognition

### v2

@@ 3 Dataset Collection and Construction
Annotating video descriptions is time-consuming and expensive. Thus, we create our dataset from a comprehensive and varied collection of videos with temporally localized chapter information, including chapter titles and start times, specifically for fitness activities. YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video. Those chapters are manually added by the uploader or automatically generated. Each chapter is a continuous, non-overlapping segment that divides the video into a video clip and its corresponding textual description.
@@
We manually collect 1388 fitness videos with pre-annotated video chapters. Additionally, we retrieve the captions automatically generated by YouTube. All our videos show a tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.). The trainers are primarily performing HIIT home workouts, focusing on bodyweight exercises with no or less equipment. The exercise videos provide instructions for the activities while showing the activities clearly. The annotation does not include predefined categories or labels. Each video owner annotates the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”. Based on the temporal boundaries of the chapters within each video, we split the videos in shorter activity clips along with their corresponding caption segments. Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter annotations. Hence, we only keep exercises shorter than 2 minutes and discard longer exercises. After that, the dataset consists of 10,695 (video, chapter, caption)-samples.
@@
We convert all 10,695 exercise videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation. To enhance the dataset with diverse modalities, we use IMUTube Kwon et al. (2021) to obtain pose and virtual sensor (accelerometer and gyroscope) features. As a result, our dataset includes (video, pose, sensor, text)-pairs for each input. For video and text we extracte multiple features. Video features include using HERO Li et al. (2020), which is utilizing either CLIP Radford et al. (2021) or S3D Xie et al. (2018) embeddings. For text, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list. Further, we use GPT-3.5 Turbo to generate detailed descriptions of the exercises. We then extract multiple text features with various models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT, as well as language features and motion tokens extracted with the Text2Motion Transformer Guo et al. (2022). We apply IMUTube to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes. We also use the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3D pose lifting. Since IMUTube fails to extract sensor data for 884 exercise videos, we exclude these sequences. We use the remaining 9811 exercises as our dataset and split it into 8,227 train and 1,584 validation activity clips. The split is based on the level of the 1,388 original fitness videos, to ensure video independence. We do not create a test set because the dataset is used solely for pre-training, not for testing. Since workout session videos typically start and end with explanations given by the trainer, we discard the initial and final 15% [^2] This value has proven to be effective in experimental investigations. of each activity clip. Despite these filtering steps, we receive a large dataset totaling around 78 hours in length. Instances are generated using sliding windows $x_{i}$ of 100 frames in length (i.e., 2 seconds) with 50 frames overlap from the videos. We found this size suitable for fitness exercises as a 2 second window generally captures enough meaningful information regarding the subject’s movements. In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, virtual IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.
@@ 4 Method
Our method leverages multimodal information for HAR pre-training. In this Section, we outline the task and its formulation, along with detailing the multimodal representation methods employing CLIP Radford et al. (2021)-like contrastive learning. Our objective is to improve HAR by employing contrastive learning across multiple modalities. The fundamental idea is that the extracted information from different modalities for a short video interval should be similar, as they represent the same activity from various perspectives. Each instance is a window, and for each one, we extract features for every modality. This means our data consists of $\mathcal{X}^{m}=\{x^{1}_{m},...,x^{N}_{m}\}$, where $x^{i}_{m}$ is the feature for modality $m$ at window $i$. We use as modalities:
  • $\mathcal{X}^{pose}$: The sequence of 3D poses for the subject extracted with ImuTube Kwon et al. (2021) utilizing OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019).
  • $\mathcal{X}^{video}$: Video features extracted with HERO Li et al. (2020) utilizing CLIP Radford et al. (2021) embeddings.
@@
Each modality $m$ has its own encoder network $e_{m}$ and projection layers $p_{m}$. Thus, $rep_{m}(x_{m})=p_{m}(e_{m}(v_{m}))$ serves as our representation for a vector $v_{m}$ of modality $m$ in our shared representation space. We learn this shared space between modality representations using the InfoNCE loss. First, for all possible $(m_{a},m_{b})$ modality pairs, we compute similarities using
@@
Then, those similarities are softmax-normalized within the batch to calculate the loss between the two modalities
@@
where $x_{a}$ and $x_{b}$ correspond to aligned batch data for modalities $a$ and $b$, respectively, and $\tau$ is the temperature hyper-parameter. Since we calculate losses between all modalities, our total loss amounts to
@@
where $M$ is the total number of modalities (in our case 4). An example can be seen in Figure 1.
Regarding the structure of our encoders, we utilize a model with a single fully connected layer for $e_{text}$ and two fully connected layers for $e_{video}$. For $e_{pose}$ and $e_{sensor}$, we employ the CNN encoder architecture as described in Strömbäck et al. (2020), making slight adaptations to accommodate our input shapes. The projection $p_{m}$ is a two-layer fully connected model for all modalities $m$, mapping them to feature vectors of length 1280.
Hyperparameter and model optimization.
We applied a random search approach for hyperparameter selection. Our hyperparameters include learning rate, number and size of projection layers, dropout rates, and activation functions. Furthermore, we experimented with various configurations of the loss function, specifically comparing the effect of using the text modality as an anchor, i.e., only considering sensor-text, pose-text, and video-text loss pairs, versus incorporating all possible modality combinations into the loss. Encompassing all modality pairs together yielded superior results. We explored weighting the text modality higher than the others in the loss, but it did not have any observable effect on performance. We also explored the use of S3D-based video embeddings, the gyroscope as sensor input, and the other text modalities and features described in Section 3 as inputs. However, since those configurations did not yield optimal performance during the hyper-parameter search, we have omitted those results for the sake of brevity.
@@ 5.1 Evaluation Datasets
We conduct experiments on the MM-Fit Strömbäck et al. (2020), MyoGym Koskimäki et al. (2017), MotionSense Malekzadeh et al. (2018), and MHEALTH Banos et al. (2015) datasets. MM-Fit is designed for multimodal fitness activity recognition and combines data from multiple modalities, including video, pose estimations, and various sensors. The dataset includes 21 workout sessions conducted by ten subjects, each performing ten different fitness exercises. The MyoGym dataset comprises 30 distinct gym exercises focused on upper body training by ten individuals, each equipped with a sensor on their right forearm. The MotionSense dataset consists of sensor data collected from 24 participants carrying an iPhone 6s in their front pocket, covering six different activities (walking downstairs, walking upstairs, sitting, standing, and jogging). The MHEALTH dataset includes twelve physical activities (such as climbing stairs, cycling, and running) performed by ten subjects, with data captured by three sensors positioned on the chest, right wrist, and left ankle. All datasets, except for MotionSense, include a "No Activity" category (NULL class) in addition to the specified activities.
@@ 5.2 Unimodal Classification on Real-World Accelerometer Data
We evaluate the performance of our pre-trained representation on real-world accelerometer data using all four datasets. For MM-Fit, we use the accelerometer data from the smartwatches worn on the left and right wrists, which we had to downsample from 100 Hz to 50 Hz via linear interpolation to match our input frequency. For MyoGym, MotionSense, and MHEALTH, we use accelerometer data from all available sensors, which are already at 50 Hz, allowing us to employ the data directly without additional preprocessing. For all four datasets, we apply $Z$-score normalization and use a sliding window of 100 samples in length (i.e., 2 seconds) with 50 samples overlap to generate instances. Furthermore, since each target dataset has different sensor placements, we have to pre-train a separate MuJo model for each one. Specifically, during pre-training our models, we use the virtual sensor positions generated by IMUTube that best match the sensor positions of the respective target dataset. Each of the four models is pre-trained on our FiMAD dataset incorporating all four input modalities: text, video, poses, and virtual accelerometer data.
@@
We conduct our experiments based on the standard experimental methodology in the HAR community, i.e., testing on unseen users and evaluating using the Macro $F_{1}$-Score, a metric that is well-suited for handling class imbalance. For MM-Fit, we use the training, validation, and cross-subject test sets as described in Strömbäck et al. (2020). Since no fixed splits are provided for MyoGym, MotionSense, and MHEALTH, we employ $k$-fold cross-validation with $k=5$. Additionally, we consider two scenarios: training the classifiers with and without the inclusion of the NULL class. This is relevant because the NULL class is not part of our pre-training dataset but is present in real-world scenarios and included in all evaluation datasets except for MotionSense.
@@
For classification, we utilize a classifier architecture $C$ comprising our sensor encoder $e_{sensor}$, projection $p_{sensor}$, and a classification head $cl$ consisting of 2 layers. We contrast three scenarios: In the baseline setting, all weights are initialized randomly, meaning the model lacks any pre-trained knowledge. In the other two scenarios, the encoder and projection head are initialized with weights from our pre-trained model. In one case, the weights of the encoder and projection head are fixed (frozen), whereas in the other case, they are trainable and we fine-tune them during the classifier’s training process. We train with different proportions of the training data, ranging from one to 100 percent, with the training data being sampled randomly per class to ensure balanced representation. To guarantee comparability of all results, we train each classifier $C$ with the same settings: Training is conducted for a maximum of 200 epochs, incorporating early stopping after 25 epochs without improvement. We employ a weighted cross-entropy classification loss to address class imbalance, given by:
$\mathcal{L}_{CL}\;\;=\;\;-\sum_{i=1}^{n}w_{y_{i}}y_{i}\log(C(x_{i}))$ (4)
@@
where $n$ is the number of classes, $y_{i}$ is a one-hot encoded ground truth activity label and $C(x_{i})$ is the predicted label vector of $C$ for input $x_{i}$, that is, the application of our $cl$ classification layers to obtain $cl(p(e(x_{i})))$. The weight $w_{y_{i}}=\frac{\max{(M)}}{m_{y_{i}}}$ for class $y_{i}$ is calculated by dividing the number of data samples of the largest class by the number of instances in class $y_{i}$ itself. To ensure validity of the results, we train 20 models independently for each configuration, with the training data being randomly selected anew for each run (for proportions less than 100%) and the non-pre-trained weights being reinitialized randomly.
5.3 Proxy task comparisons
@@
We compare our approach with Multi-task Saeed et al. (2019), Auto-encoder Haresamudram et al. (2019) and SimClr Tang et al. (2021). All approaches were trained with the same simulated IMU data as our approach and use the same encoder architecture as ours, for fair comparison. Regarding other configurations SimClr used random rotations as augmentation uniformely sampled in $(-\pi,\pi)$ range. The representation is obtained from our projection layer through two dense layers of size 256, 128 and 128, respectively, with the first two being followed by a relu activation. It was trained with the NTXent loss using l2 norm. For Multitask we have selected as augmentations noising, scaling, negation, horizontally flipping, permutation and channel shuffling. Noising consists of adding a value sampled from a normal distribution with mean zero and standard deviation of 0.05 while scaling consists of multiplying by values drawn from this same distribution. Negation, horizontally flipping and channel shuffling are self evident. Permutation means randomly splitting the time series into four parts and changing their order. Every task has a head that takes as input the encoder’s output and consists of one fully connected layer with 256 size and relu activation, followed by another fully connected layer with also 256 size but sigmoid activation. For Auto-encoder the representation is passed to a convolutional decoder that re-creates the original signal. This decoder consists of first a linear layer to bring the projection to $(32,n\_sensors,\frac{win\_size}{4})$ where $n\_sensors$ is the number of sensors in the dataset and $win\_size$ is the original window size. This is followed by batch norm, relu and then three blocks of convolution transposed, $0.05$ dropout, relu and upsampling. A kernel size of $(1,11)$ was used and the channels change following the sequence $(32,16,8,orig)$, with $orig$ being the original number of sensor channels. To our knowledge, this is the first application of such methods to simulated sensor data, as well as their first use in a multi-sensor scenario. While all those methods could be added to our framework easily, we selected them as comparison to highlight how our method can learn from extra modalities only available at training time beyond what is present at the simulated sensor data itself (which is the only input for those methods).

### v3

@@ 3 Dataset Collection and Construction
Annotating video descriptions is time-consuming and expensive. Thus, we created our dataset from a comprehensive and varied collection of videos with temporally localized chapter information, including chapter titles and start times, specifically for fitness activities. YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video. Those chapters are manually added by the uploader or automatically generated. Each chapter is a continuous, non-overlapping segment that divides the video into a video clip and its corresponding textual description.
@@
We manually collected 1388 fitness videos with pre-annotated video chapters. Additionally, we retrieved the captions automatically generated by YouTube. All our videos show a tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.). The trainers primarily conduct HIIT home workouts, focusing on bodyweight exercises with little to no equipment. The exercise videos provide instructions for the activities while demonstrating them clearly. The annotation does not include predefined categories or labels as each video owner annotated the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”. Based on the temporal boundaries of the chapters within each video, we split the videos into shorter activity clips along with their corresponding caption segments. Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter annotations. Hence, we only kept exercises that are shorter than 2 minutes and discarded longer exercises. After that, the dataset consists of 10,695 (video, chapter, caption)-samples.
@@
We converted all 10,695 exercise videos to 720p resolution and 50 frames per second via FFmpeg’s adaptive overlapped block motion compensation with bidirectional motion estimation. To enhance the dataset with diverse modalities, we used IMUTube (Kwon et al., 2021a) to obtain poses and virtual sensor (accelerometer and gyroscope) features. As a result, our dataset includes (video, pose, sensor, text)-pairs for each input. For video and text, we extracted multiple features. The video features were generated using HERO (Li et al., 2020), using CLIP (Radford et al., 2021) and S3D (Xie et al., 2018) embeddings. For the text modality, we applied several normalization strategies to the raw chapters, such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and some special symbols from a predefined list. In addition, we used GPT-3.5 Turbo to generate detailed descriptions of the exercises. We then extracted multiple text features with various models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT, as well as language features and motion tokens extracted with the Text2Motion Transformer (Guo et al., 2022). We applied IMUTube to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes. We also used the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3D pose lifting. Since IMUTube failed to extract sensor data for 884 exercise videos, we excluded these sequences. We used the remaining 9,811 exercises as our dataset and split it into 8,227 train and 1,584 validation activity clips. The split is based on the level of the 1,388 original fitness videos to ensure video independence. We did not create a test set because the dataset is used solely for pre-training, not for testing. Since workout session videos typically start and end with explanations given by the trainer, we discarded the initial and final 15% [^2] This value has proven to be effective in experimental investigations. of each activity clip. Despite these filtering steps, we receive a large dataset totaling around 78 hours in length. Instances were generated from the videos using sliding windows $x_{i}$ of 100 frames in length (i.e., two seconds) with an overlap of 50 frames. We found this size suitable for fitness exercises, as a two-second window generally captures enough meaningful information regarding the subject’s movements.
In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, virtual IMU data, chapter labels, captions, and GPT-generated descriptions, as well as several types of features extracted from these text modalities.
@@ 4 Method
Our method leverages multimodal information for HAR pre-training. In this section, we outline the task and its formulation, along with detailing the multimodal representation methods employing CLIP (Radford et al., 2021)-like contrastive learning. Our objective is to improve HAR by employing contrastive learning across multiple modalities. The fundamental idea is that the information extracted from different modalities for a short video interval should be similar, as they represent the same activity from various perspectives.
Each instance is represented as a window, and for each window, we extract features for every modality. This means that our data consists of $\mathcal{X}^{m}=\{x^{1}_{m},...,x^{N}_{m}\}$, where $x^{i}_{m}$ is the feature of modality $m$ in window $i$. The modalities we utilize include:
  • $\mathcal{X}^{pose}$: The sequence of 3D poses for the subject extracted with ImuTube (Kwon et al., 2021a) utilizing OpenPose (Cao et al., 2017) and VideoPose3D (Pavllo et al., 2019).
  • $\mathcal{X}^{video}$: Video features extracted with HERO (Li et al., 2020) using CLIP (Radford et al., 2021) embeddings.
@@
Each modality $m$ has its own encoder network $e_{m}$ and projection layers $p_{m}$. Thus, $rep_{m}(x_{m})=p_{m}(e_{m}(v_{m}))$ serves as our representation for a vector $v_{m}$ of modality $m$ in our shared representation space. We learn this shared space between the modality representations using the InfoNCE loss. First, for all possible sets of two modalities $\{m_{a},m_{b}\}$, we compute similarities using
@@
Then, those similarities are softmax-normalized within the batch to calculate the loss between the two modalities.
@@
where $x_{a}$ and $x_{b}$ correspond to aligned batch data for modalities $a$ and $b$, respectively, and $\tau$ is the temperature hyper-parameter  [^3] Best results obtained with $\tau=1.0$.. In other words, for each data instance, its representations across different modalities (positive) should be more similar to each other than at other instances (negatives, which are all other examples in the same batch). As in CLIP, we compute $\mathcal{L}_{c}(x_{m_{a}},x_{m_{b}})+\mathcal{L}_{c}(x_{m_{b}},x_{m_{a}})$ for each pair, symmetrically aligning the embeddings and covering the negatives from both perspectives. Since we calculate losses between all modalities, our total loss amounts to
@@
where $M$ is the total number of modalities. In our case, four modalities generate six combinations, as shown in Figure 1.
Hyperparameter search and model optimization
We applied a random search approach to identify the optimal hyperparameters for our model. Our hyperparameters included the number and size of encoder and projection layers, learning rate, dropout rates, and activation functions. Furthermore, we experimented with various configurations of the loss function, specifically comparing the effect of using the text modality as an anchor, i.e., only considering sensor-text, pose-text, and video-text loss pairs, versus incorporating all possible modality combinations into the loss. Encompassing all modality pairs together yielded superior results. We explored weighting the text modality higher than the others in the loss, but it did not have any observable effect on performance. We also explored the use of S3D-based video embeddings, the gyroscope as the sensor input, and the other text modalities and features described in Section 3 as inputs. However, since those configurations did not yield optimal performance during the hyperparameter search, we have omitted those results for the sake of brevity.
Model structure
Based on the results of the hyperparameter tuning, we designed our encoder structures to include a model with a single fully connected layer for $e_{text}$ and two fully connected layers for $e_{video}$. $e_{text}$ has an input size of 1536 (the size of the ADA text embeddings) and an output size of 768. $e_{video}$ has an input size of 1024 (size of the HERO CLIP features), a hidden layer size of 256, and an output size of 256. For $e_{pose}$ and $e_{sensor}$, we employed the CNN encoder architecture as described in (Strömbäck et al., 2020), making slight adaptations to accommodate our input shapes. Both encoders consist of three CNN layers. $e_{pose}$ has a kernel size of (11, 11) and includes two fully connected layers, each with a size of 1024. $e_{sensor}$ has a kernel size of ($s_{n}$, 11) for the first layer, where $s_{n}$ represents the number of sensor inputs. The kernel sizes for layer 2 and layer 3 are (1, 11), followed by a single fully connected layer with an output size of 512. The projections $p_{m}$ are fully connected three-layer models for all modalities $m$, with hidden and output layer sizes of 1280. All encoders and projections utilize GELU as activation function and dropouts of 0.4 for the projections, 0.3 for $e_{video}$, 0.6 for $e_{text}$, 0.4 for $e_{pose}$, and no dropout for $e_{sensor}$. We trained with a batch size of 256 for a maximum of 200 epochs with early stopping (patience = 50). We utilized the AdamW optimizer with a learning rate of 0.0008 in combination with a cosine annealing scheduling for the learning rate during training that starts with 34 warm up epochs followed by a maximum number of 33 iterations using a minimum learning rate of 3.0398e-06.
@@ 5.1 Evaluation Datasets
We conducted experiments on the MM-Fit (Strömbäck et al., 2020), MyoGym (Koskimäki et al., 2017), MotionSense (Malekzadeh et al., 2018), and MHEALTH (Banos et al., 2015) datasets. MM-Fit is designed for multimodal recognition of fitness activities and combines data from multiple modalities, including video, pose estimations, and various sensors. The dataset includes 21 workout sessions conducted by ten subjects, each performing ten different fitness exercises. The MyoGym dataset comprises 30 distinct gym exercises focused on upper body training by ten individuals, each equipped with a sensor on their right forearm. MotionSense consists of sensor data collected from 24 participants carrying an iPhone 6s in their front pocket, covering six different activities (walking downstairs, walking upstairs, sitting, standing, and jogging). The MHEALTH dataset includes twelve physical activities (such as climbing stairs, cycling, and running) performed by ten subjects, with data captured by three sensors positioned on the chest, right wrist, and left ankle. All datasets, except for MotionSense, include a “No Activity” category (NULL class) in addition to the specified activities.
@@ 5.2 Unimodal Classification on Real-World Accelerometer Data
We evaluated the performance of our pre-trained representation on real-world accelerometer data using all four datasets. For MM-Fit, we used the accelerometer data from the smartwatches worn on the left and right wrists, which we had to downsample from 100 Hz to 50 Hz via linear interpolation to match our input frequency. For MyoGym, MotionSense, and MHEALTH, we used accelerometer data from all available sensors, which are already at 50 Hz, allowing us to employ the data directly without additional preprocessing. For all four datasets, we applied $Z$-score normalization and used a sliding window of 100 samples in length (i.e., two seconds) with 50 samples overlap to generate instances. Furthermore, since each target dataset has different sensor placements, we had to pre-train a separate MuJo model for each one. Specifically, during pre-training our models, we used the virtual sensor positions generated by IMUTube that best match the sensor positions of the respective target dataset. Each of the four models was pre-trained on our FiMAD dataset incorporating all four input modalities: text, video, poses, and virtual accelerometer data.
@@
We conducted our experiments based on the standard experimental methodology in the HAR community, i.e., testing on unseen users and evaluating using the Macro $F_{1}$-Score, a metric that is well-suited for handling class imbalance. For MM-Fit, we used the training, validation, and cross-subject test sets as described in (Strömbäck et al., 2020). Since no fixed splits are provided for MyoGym, MotionSense, and MHEALTH, we created five splits, where we randomly divided subjects into independent training, validation, and test sets for each split. Additionally, we considered two scenarios: training the classifiers with and without the inclusion of the NULL class. This is relevant because the NULL class is not part of our pre-training dataset but is present in real-world scenarios and is included in all evaluation datasets except for MotionSense.
@@
For classification, we utilized a classifier architecture $C$ comprising our sensor encoder $e_{sensor}$, projection $p_{sensor}$, and a classification head $cl$ consisting of two layers. We contrast three scenarios: In the baseline setting, all weights were initialized randomly, meaning the model lacks any pre-trained knowledge. In the other two scenarios, the encoder and projection head were initialized with weights from our pre-trained model. In one case, the weights of the encoder and projection head were fixed (frozen), whereas in the other case, they were trainable and we fine-tuned them during the classifier’s training process. We trained with different proportions of the training data, ranging from one to 100 percent, with the training data being sampled randomly per class to ensure balanced representation. To guarantee the comparability of all results, we trained each classifier $C$ with the same settings: Training was conducted for a maximum of 200 epochs, incorporating early stopping after 25 epochs without improvement. We employed a weighted cross-entropy classification loss to address class imbalance, given by
$\mathcal{L}_{CL}\;\;=\;\;-\sum_{i=1}^{n}w_{y_{i}}y_{i}\log(C(x_{i})),$ (4)
@@
where $n$ is the number of classes, $y_{i}$ is a one-hot encoded ground truth activity label and $C(x_{i})$ is the predicted label vector of $C$ for input $x_{i}$, that is, the application of our $cl$ classification layers to obtain $cl(p(e(x_{i})))$. The weight $w_{y_{i}}=\frac{\max{(M)}}{m_{y_{i}}}$ for class $y_{i}$ is calculated by dividing the number of data samples of the largest class by the number of instances in class $y_{i}$ itself. To ensure the validity of the results, we trained 20 models independently for each configuration, with the training data being randomly selected anew for each run (for proportions less than 100%) and the non-pre-trained weights being reinitialized randomly.
Proxy task comparisons
@@
We compared our approach with Multi-task (Saeed et al., 2019), Auto-encoder (Haresamudram et al., 2019), and simCLR (Tang et al., 2021). For a fair comparison, all approaches were trained with the same simulated IMU data and the same encoder architecture as our approach. Regarding other configurations, simCLR uses random rotations as augmentation uniformly sampled in $(-\pi,\pi)$ range. The representation was obtained from our projection layer through two dense layers of size 256, 128, and 128, respectively, the first two being followed by a ReLU activation. It was trained with the NTXent loss using the l2 norm. For Multi-task, we have selected the following augmentations: noising, scaling, negation, horizontal flipping, permutation, and channel shuffling. Noising consists of adding a value sampled from a normal distribution with a mean of 0.0 and a standard deviation of 0.05, while scaling consists of multiplying by values drawn from this same distribution. Negation, horizontal flipping, and channel shuffling are self-evident. Permutation means randomly splitting the time series into four parts and changing their order. Every task has a head that takes as input the encoder’s output and consists of one fully connected layer with 256 size and ReLU activation, followed by another fully connected layer with also 256 size but sigmoid activation. For Auto-encoder, the representation is passed to a convolutional decoder that re-creates the original signal. This decoder consists of first a linear layer to bring the projection to $(32,n\_sensors,\frac{win\_size}{4})$ where $n\_sensors$ is the number of sensors in the dataset and $win\_size$ is the original window size. This is followed by batch norm, ReLU and then three blocks of convolution transposed, $0.05$ dropout, ReLU and upsampling. A kernel size of $(1,11)$ was used and the channels change following the sequence $(32,16,8,orig)$, with $orig$ being the original number of sensor channels.
To our knowledge, this is the first application of such methods to simulated sensor data, as well as their first use in a multi-sensor scenario. While all those methods could be easily added to our framework, we selected them as a comparison to highlight how our method can learn from extra modalities that are only available during training time, beyond what is present in the simulated sensor data itself (which is the only input for those methods).